# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@
Our results further point to a modular device with distinct quantum processing and sensing modules, realizing a vision of a quantum computer connected to a high-precision sensor (Fig. 5e). The processing module could be used to prepare entangled probe states in parallel to sensing, followed by a state exchange operation between the sensor and the processor. Subsequently, readout could be performed in the processing module in parallel to sensing, before preparing new entangled probe states, yielding very low dead time operation. Options for fully realizing this vision could be based on a multi-zone approach with a single species 24, a dual-species 56 device composed of alkaline-earth and alkali atoms (e.g., strontium and rubidium), an optical-metastable-ground-state qubit architecture 59, 64 or combinations thereof. Several of our results can be viewed as proof-of-principle demonstrations towards this modular vision, including state exchange operations (Ext. Data Fig. 11) and higher-weight ancilla-based readout. Such a device could be part of next-generation quantum technologies from field-transportable sensors to state-of-the-art optical clocks 35 in the quest to push the limits of precision measurements.
@@ Acknowledgements
Data and code availability
The data and codes supporting this study’s findings are available from the corresponding author upon reasonable request.

### v3

@@
Our results further point to a modular device with distinct quantum processing and sensing modules, realizing a vision of a quantum computer connected to a high-precision sensor (Fig. 5e). The processing module could be used to prepare entangled probe states in parallel to sensing, followed by a state exchange operation between the sensor and the processor. Subsequently, readout could be performed in the processing module in parallel to sensing, before preparing new entangled probe states, yielding very low dead time operation. Options for fully realizing this vision could be based on a multi-zone approach with a single species 24, a dual-species 40 device composed of alkaline-earth and alkali atoms (e.g., strontium and rubidium), an optical-metastable-ground-state qubit architecture 56, 60 or combinations thereof. Several of our results can be viewed as proof-of-principle demonstrations towards this modular vision, including state exchange operations (Ext. Data Fig. 11) and higher-weight ancilla-based readout. Such a device could be part of next-generation quantum technologies from field-transportable sensors to state-of-the-art optical clocks 35 in the quest to push the limits of precision measurements.
@@ Acknowledgements
Data availability
The data supporting this study’s findings are available from the corresponding author upon reasonable request.
Code availability
The codes supporting this study’s findings are available from the corresponding author upon reasonable request.